The Average mass of 50 pumpkin is 2,1kg. If one more pumpkin is added, the average mass is 2,2kg. What is the mass of the extra pumpkin?


1
Expert's answer
2021-07-19T05:48:14-0400

"\frac{2.1*50+x}{51}=2.2."

"105+x=112.2."

"x=112.2-105=7.2kg."